Steel railroad rails 18,2880m long are laid on a winter day when the temperature is - 6,00 degrees Celsius. How much space must be left between rails if they are just to touch on a summer day when the temperature is 43 degree Celsius.
1
Expert's answer
2020-08-02T15:04:23-0400
"\Delta l=l_0\alpha \Delta T\\\Delta l=(18.288)(1.2\cdot10^{-5})(43+6)=0.0108\ m"